OpenStudy (anonymous):

Can you check these two questions, please? What is the product of 2 sqrt(u( sqrt(5)+ 3 sqrt(45))? Simplify if possible. A) 20 sqrt(5)u B) 20 sqrt(5)u + 9 sqrt(5) C) 56 sqrt(5)u D) 90u sqrt(5) I think it's A.20 sqrt(5)u Which expression is the simplified form of sqrt(k^21)? A) k^5 B) k^10 sqrt(k) C) k^10 D) k^4 sqrt(k) Again, I think it's A. k^5

OpenStudy (phi):

k^5 * k^5 = k^10 in other words sqrt(k^10) = k^5 so it does not make sense that sqrt(k^21) also equals k^5